OpenStudy (anonymous):

Simplify : x − 4y + 2z − 3x − 3y + 2z = ?

OpenStudy (anonymous):

Combine like terms. \[(x - 3x) + (-4y - 3y) + (2z + 2z) = ?\]

OpenStudy (anonymous):

-2x2-7y2+4z

OpenStudy (anonymous):

@doubledrive How in the world did you get the squared parts? There is no multiplying, just addition and subtraction.

OpenStudy (anonymous):

lol, doesnt the x and y get squared since there are two??

OpenStudy (anonymous):

Nope. You're adding the like terms. x - 3x = -2x -4y - 3y = -7y 2z + 2z = 4z Wait, if you added the number of x and y, then how come you didn't do the same for z? It just doesn't make sense lol :)

OpenStudy (anonymous):

so whats the answr?

OpenStudy (anonymous):

oh, the z, well i forgot about that letter